OpenStudy (anonymous):

Use L'Hospitals Rule lim theta rightarrow 0 \frac{ theta }{ tan(theta) }

OpenStudy (anonymous):

I know that we are suppose to substitute the zero but with the trig functions, I am not sure what to do

OpenStudy (anonymous):

i got 0/tan 0

OpenStudy (anonymous):

= zero and tan 0 is just zero right?

OpenStudy (anonymous):

l'hopital rule states that both denominator and numerator can be derivated when looking for lim if lim of both original functions is either 0 or infinite limes of both theta and tan(theta) when theta approaches 0 are 0 now we derivate both theta and tan(theta) \[\lim_{\theta \rightarrow 0} \frac{ \theta \prime }{ \tan(\theta)\prime } = \frac{ 1 }{ \frac{ 1 }{ \sin ^{2} \theta } }\] can you find the limes now?

OpenStudy (anonymous):

1/sin^2theta = sin^2theta

OpenStudy (anonymous):

do i take the reciprocal of sin^2 theta

OpenStudy (anonymous):

sorry, my bad. i derivated tangent wrong. tan(theta) derivated is 1/cos^2 theta

OpenStudy (anonymous):

ok- I dont know why i thought it was sin and cos for tan

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0} \frac{ \theta \prime }{ \tan(\theta)\prime } = \frac{ 1 }{ \frac{ 1 }{ \cos ^{2} \theta } }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then you have: \[=\frac{ \cos ^{2}\theta }{ 1 }\]

OpenStudy (anonymous):

do i take the recipricol of that?

OpenStudy (anonymous):

nope, you just plug in the value of theta as is defined in the limes and see what you get

OpenStudy (anonymous):

so cos^2 (0) = 1 right?

OpenStudy (anonymous):

L'Hopital rule is used when you have limes of a fraction (or something that can be rewritten as fraction) but both numerator and denominator approach either 0 or infinite

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok thank you so much for your help

OpenStudy (anonymous):

np;) here is some extra reading if you need: http://en.wikipedia.org/wiki/L%27H%C3%B4pital%27s_rule

OpenStudy (anonymous):

awsome, thanks

OpenStudy (anonymous):

You can also use the fact that \(\tan\theta\approx\theta\) for \(\theta\) near 0, giving you \[\lim_{\theta\to0}\frac{\theta}{\theta}\] Which you can evaluate using l'Hopital's rule, or simplify immediately.